OpenStudy (idealist10):

Find a particular solution of y"+3y'+2y=7cosx-sinx.

OpenStudy (idealist10):

@Compassionate @sammixboo @thomaster @zepdrix

zepdrix (zepdrix):

Particular? It will have the "form" of the right side, ya?\[\large\rm y_p=A \cos x+ B \sin x\]

zepdrix (zepdrix):

\[\large\rm y_p'=-A \sin x+ B \cos x\]\[\large\rm y_p''=-A \cos x- B \sin x\]

zepdrix (zepdrix):

And then ummm, substitute that stuff back into the original equation. Ya? :o

OpenStudy (idealist10):

Let me try.

zepdrix (zepdrix):

Recommendation: Call C=cos(x) and S=sin(x) for problems like this. It's so tedious to rewrite sine and cosine over and over lol

OpenStudy (idealist10):

So I got Acosx+3Bcosx=7cosx and Bsinx-3Asinx=-sinx, how to find A and B?

zepdrix (zepdrix):

\[\large\rm A \cos x+3B \cos x=7\cos x\]Factor a cosine out of each term,\[\large\rm (A+3B)\cos x=7\cos x\]Divide out the cosines,\[\large\rm A+3B=7\]

zepdrix (zepdrix):

Understand those steps? Try it with the sines :)

OpenStudy (idealist10):

I was stupid. Thanks.